40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3859     Entry time: Thu Nov 4 03:13:46 2010
Author: Suresh 
Type: Update 
Category: Locking 
Subject: Fibre coupling 1064nm light at the south-end table 

[Kiwamu, Suresh]

We decided to use the 1064nm beam reflected from the Y1-1037-45-P mirror after the collimation lens following the doubling crystal for coupling into the optical fiber (ref 3843 and 3847 ).

We replaced a beam dump which was blocking this beam with a Y1-1037-45-P mirror and directed the beam into the fiber coupler with another Y1-1037-45-P.  The power in this beam was about 1W.  This has been stepped down to 10mW by introducing a reflective ND filter of OD=2.  The reflected power has been dumped into a blade-stack beam dump.

Steve has ordered the The Visual Fault Locator from Fluke.  It is expected to arrive within a day or two.

 

 

ELOG V3.1.3-